a constant force acts on a 5.0kg object and reduces its velocity from 7.0 m/s to 3.0 m/s in a time of 3.0 s. Find the force


1
Expert's answer
2021-09-02T10:41:30-0400

Given:

m=5.0kgm=5.0\:\rm kg

vi=7.0m/sv_i=7.0\:\rm m/s

vf=3.0m/sv_f=3.0\:\rm m/s

Δt=3.0s\Delta t=3.0\:\rm s


The acceleration of an object

a=ΔvΔt=3.07.03.0=1.3m/s2a=\frac{\Delta v}{\Delta t}=\frac{3.0-7.0}{3.0}=-1.3\:\rm m/s^2

The magnitude of force

F=ma=5.01.3=6.7NF=ma=5.0*1.3=6.7\:\rm N